OpenStudy (anonymous):

Chargaff's rules of DNA structure states that A. a+t=g+c B.a+g=t+c C. a=c,t=g D. a=g, t=c E. the number of purines in DNA never match the number of pyrimidines i know it's not e,b, or d

OpenStudy (anonymous):

a+g = t+c

OpenStudy (anonymous):

B

OpenStudy (anonymous):

why is it b

OpenStudy (anonymous):

because in b, the ratio of purines to pyrimidines is 1:1

OpenStudy (anonymous):

since a pairs with t and g pairs with c, you will get 50% for each side

OpenStudy (anonymous):

ah i see cuz the book says a=t and g=c

OpenStudy (anonymous):

yeah that's right. but a-t =/= g-c

OpenStudy (anonymous):

that's also right. a=t and g=c because those bases pair with each other.